Q 20) A kittle holds 2.00 kg of water at 10°C. Energy required to heat the water to the boiling point at

sea level ( 100°C) equals: (specific heat capacity cw = 4186 J/(kg.K):

1) 75.3 kJ

2) 753 kJ

3) 7.53 kJ

4) 75.3 kJ


1
Expert's answer
2021-05-16T17:52:26-0400

Find this energy:


"Q=cm(T_2-T_1)=4186·2(100-10)=\\=753.48·10^3\text{ kJ}."